Compute the amount of each foreign currency that can be purchased for $2 million:

  1. 1 Korean won = $0.0004 (Round your answer to the nearest whole number.)

  1. 1 Malaysian ringgit = $0.3244 (Round your answer to the nearest whole number.)

  1. 1 Thai baht = $0.0337 (Round your answer to the nearest whole number.)
Answer #1
a) Korean won = 2000000/0.0004 = 5000000000 WON
b) ringgit = 2000000/0.3244 = 6165228 ringgit
c) Thai baht = 2000000/0.0337 = 59347181 Thai baht